OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). (1 points) Answer -9x + 27 -9x + 25 -9x + 5 -9x - 2

OpenStudy (hba):

Put g(x)=-x+3 in f(x) in place of x Got it ?

OpenStudy (hba):

@InsanelyChaotic Respond .

OpenStudy (anonymous):

wow, ok geez

OpenStudy (anonymous):

I got a

OpenStudy (anonymous):

9 x 3= 27

OpenStudy (anonymous):

@hba

OpenStudy (hba):

Do it again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

c

OpenStudy (anonymous):

I got 5!!

OpenStudy (hba):

Noooooooooooo

OpenStudy (hba):

f(x)= 9x - 2 f(g(x) or f(-x+3) f(g(x)=9(-x+3)-2 = Now solve

OpenStudy (hba):

@InsanelyChaotic Why you no understand ?

OpenStudy (anonymous):

-2

OpenStudy (hba):

^ Whaaaaaat ? :o

OpenStudy (hba):

Which option ?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@hba

OpenStudy (hba):

No

OpenStudy (hba):

Now you are left with only one 9(-x+3)-2 -9x+27-2 -9x+25 Therfore it is B

OpenStudy (anonymous):

thank you sooooo sooo soooo much

OpenStudy (hba):

yo need to concentate :/ Welcome Welcome :)

OpenStudy (anonymous):

i try it's just that math is really hard for me. It's hard to focus on something that barely makes sense to you.

OpenStudy (hba):

You know what ? When you go to a shop you spend 7$ and you take that 10 $ note and you get back 3 $ That is all maths ya see. So it makes sorta sense. You need to practice maths to be perfect. So try to practice rather than just look up for answers because if ya practice well you are not gonna have to ask for help in the future. Rather you would be teaching others.

OpenStudy (anonymous):

haha i do like to buy shirts.... good example

OpenStudy (anonymous):

When i practice, I do need the answers afterwards

OpenStudy (anonymous):

a lot of this is to practice for the big tests i do on my own

OpenStudy (hba):

So if ya dont know the maths you are gonna go in loss :P In your school life and in your real life when you get that 70 % off :P

OpenStudy (anonymous):

another thing is, this kind of math won't be in my future. every adult i talk to say that none of this is incorporated in their everyday living

OpenStudy (anonymous):

hahaha ;)))

OpenStudy (anonymous):

But i'm serious. I'm not going to need to memorize every formula. the profession I'm getting into requires none of this at all. It's pointless in my eyes

OpenStudy (hba):

Well it is sorta there in everyone life. You will get to know the use later. First learn how to use the tool then we will tell you where to work it through.

OpenStudy (anonymous):

Basic math, yes. Of course I need to know basic math, but as for all of this other material i'm learning, I won't need to refer back to this

OpenStudy (hba):

Whatcha gonna become ?

OpenStudy (anonymous):

A professional makeup artist

OpenStudy (anonymous):

slash stripper

OpenStudy (anonymous):

hahaha just kidding about the second one

OpenStudy (hba):

Yeah right :P But ya need to do the maths so if you do not get a job in that part you can earn money by teaching some kids at a local school :P

OpenStudy (hba):

Education(Learning) never goes useless.

OpenStudy (anonymous):

Trust me, I'm becoming a makeup artist

OpenStudy (anonymous):

If not that, a paralegal

OpenStudy (hba):

Trust me,There me education never goes usless.

OpenStudy (anonymous):

or a Hooter's girl for some time

OpenStudy (hba):

lol your kids gonna be like,mom,you ain't know anything.

OpenStudy (anonymous):

I know it never goes useless :) it seems like it though, sometimes

OpenStudy (hba):

The sometimes part is for the future then. You will get to know its use one day.

OpenStudy (anonymous):

thank you

OpenStudy (hba):

Welcome.